Burklin, Inc., has earnings of $18.5 million and is projected to grow at a constant rate of 5 percent forever because of the benefits gained from the learning curve. Currently, all earnings are paid out as dividends. The company plans to launch a new project two years from now that would be completely internally funded and require 25 percent of the earnings that year. The project would start generating revenues one year after the launch of the project and the earnings from the new project in any year are estimated to be constant at $7 million. The company has 8.3 million shares of stock outstanding.

    

Estimate the value of the stock. The discount rate is 10 percent.
Answer #1

Ans. Value of Stock = $48.4

Dividend for first 2 years = Net earnings

Dividend per share = Dividend / no. of shares = 18.5 / 8.3 = $2.229

Dividend in year 1 = 2.229 * (1+5%) = $2.34

Dividend in year 2 = 2.34 * (1+5%) = $2.457

Dividend in year 3 (25% of earnings were used for the project)= 2.457 * (1+5%) * (1-25%) = $1.935

Dividend in year 4 (excl. net income from new project) = 2.457 * (1+5%) * (1+5%) = $2.709

Net income from new project constant forever = $7 million

Dividend per share from the new project (will be constant forever) = 7 / 8.3 = $0.843

Value of stock = Dividend in year 1 / (1+10%)^1 + Dividend in year 2 / (1+10%)^2 + Dividend in year 3 / (1+10%)^3+ (Dividend in year 4 (excl. net income from new project) / (discount rate - dividend growth rate)) / (1+10%)^4 + (Dividend per share from the new project / discount rate) / (1+10%)^4

Value of stock = (2.34/1.1) + (2.457/1.1^2)+ (1.935/1.1^3)+((2.709 / (10% - 5%)) / 1.1^4)+ ((0.843 / 10%)/ (1+10%)^4)

Value of stock = 2.127 + 2.031 + 1.454 + 37.006 + 5.758

Value of stock = $48.4